A teacher has two math classes. In one class there are 80 students and the average grade on a test was 30. In her other class there are 40 students and the average grade on the same test is 50. What is the average grade of all her students?
Found 2 solutions by dfrazzetto, stanbon:
Answer by dfrazzetto(283)

To find the cumulative (total) average of all students, we simply multiply each class average by the number of students, add the 2 sums together, and divide by the total number of students.
total students = 80 + 40 = 120 students
cumulative average = sum(class1 and class 2) / 120 students
= (80x30 + 40x50) / 120
= (2400 + 2000) / 120
= (4400) / 120
=36.6666666666666666 = 36 2/3

Answer by stanbon(75887)

A teacher has two math classes.
In one class there are 80 students and the average grade on a test was 30.
Total points = 80*30 = 2400
--------------------
In her other class there are 40 students and the average grade on the same test is 50.
Total points = 40*50 = 2000
--------------------
What is the average grade of all her students?
Aver = total points/total # of students
= (2400+2000)/(80+40)
= 4400/120
= 36 2/3
===============
Cheers,
Stan H.
===============